There are three candidates for the position of a head teacher-Ochieng, Abdul and Njeri, whose probabilities of getting the appointment are and  respectively. The probability that Ochieng if selected will introduce co-education in the school is 0.3.The probability of Abdul and Njeri doing the same are respectively 0.5 and 0.8.What is the probability that there will be co-education in school?                                                    


1
Expert's answer
2021-10-14T13:05:33-0400

Let "A" denote the event that Abdul  will be seleted.

Let "B" denote the event that Njeri  will be seleted.

Let "C" denote the event that Ochieng  will be seleted.

Let "E" denote the event that co-education will be in school.

Given "P(A)=2\/9, P(B)=1\/3, P(C)=4\/9,"

"P(E|A)=0.5, P(E|B)=0.8, P(E|C)=0.3."

By the Law of Total Probability


"P(E)=P(A)P(E|A)+P(B)P(E|B)+P(C)P(E|C)"

"=\dfrac{2}{9}(0.5)+\dfrac{1}{3}(0.8)+\dfrac{4}{9}(0.3)=\dfrac{4.6}{9}=\dfrac{23}{45}"

The probability that there will be co-education in school is "\dfrac{23}{45}."